OpenStudy (anonymous):

partial fraction decomposition (a s^2+b s+c)/(s^3 (s+d)) where a,b,c,d are known constants

hartnn (hartnn):

hi, do you know normal partial process, with say (Ax+B)(Cx+D) type of denominators ?

OpenStudy (anonymous):

yes , I just havent' done partials in a very long time and could use some help

OpenStudy (anonymous):

i can solve it, just determining numerators and denominators is tripping me up, can you help?

hartnn (hartnn):

sure

hartnn (hartnn):

\(\Large \dfrac{as^2+bs+c}{s^3(s+d)}=\dfrac{A}{s+d}+\dfrac{B}{s}+\dfrac{C}{s^2}+\dfrac{D}{s^3}\) do you understand this form ?

hartnn (hartnn):

and have any idea how to get A,B,C,D ??

hartnn (hartnn):

still there ? bdw, even i love to laugh :D

OpenStudy (anonymous):

no they're just known constants

hartnn (hartnn):

ok

hartnn (hartnn):

do you know how to take the LCM while adding fractions ?

hartnn (hartnn):

like for example x/y +z/y^2 will be (xy+z)/y^2

hartnn (hartnn):

from this , \(\Large \dfrac{as^2+bs+c}{s^3(s+d)}=\dfrac{A}{s+d}+\dfrac{B}{s}+\dfrac{C}{s^2}+\dfrac{D}{s^3}\) could you understand how i get \(as^2+bs+c=A(s^3)+Bs^2(s+d)+Cs(s+d)+D(s+d)\)

hartnn (hartnn):

from this, its not that difficult, to get A put s=-d